OpenStudy (anonymous):

-9 -(-3 - 5t)= -3(7t+4)

OpenStudy (anonymous):

Solving for t?

OpenStudy (jadeishere):

-9 +3 + 5t = -21t - 12 So do you know the rule that if you multiply a negative by a negative it turns positive?

OpenStudy (anonymous):

I do know that rule and just about anyway I do this problem and check my answer it's wrong lol. I believe it is -9 -(-3). I'm just to the point where I am confusing myself worse.

OpenStudy (anonymous):

but yes solving for T.

OpenStudy (anonymous):

When there is a parenthesis between a minus and negative sign it becomes a +

OpenStudy (anonymous):

Basically what I keep getting is; -9 - (-3 - 5t)= -3(7t + 4) 6-5t =-3(7t+4) 6-5t = -21t +12 6+16t= +12 16t =6 then the division is 16/6 but I get 2.666666667 and plugging that back in it doesnt equal the other side...